What volume of a 1.6 N H2SO4 solution is required to completely neutralize 50 mL of a 1.3 N NaOH?
1
Expert's answer
2020-12-01T09:18:20-0500

Volume of 1.6 N H2SO4 solution = x mL

Then 1.6*x = 1.3*50 because H2SO4 solution needs to have the same amount of H+ as there are OH- in NaOH solution to achieve neutralization

Volume of 1.6N H2SO4 required = 50*1.3/1.6 = 40.625 mL